OpenStudy (anonymous):

Simplify 16 divided by 2 + 6 – 4 x 3

hero (hero):

Is it: 16÷2+6−4 ×3

hero (hero):

16÷2+6−4 ×3 =8 + 6 - 12 =14 - 12 =2

OpenStudy (anonymous):

thanks broo

OpenStudy (anonymous):

If it's 16/(2+6-4*3) then: 16/(8-12)= =16/(-4)= =-4

OpenStudy (radar):

This is an example where the question is questionable .lol